### v1

@@ Abstract
Despite its clinical utility, medical image segmentation (MIS) remains a daunting task due to images’ inherent complexity and variability. Vision transformers (ViTs) have recently emerged as a promising solution to improve MIS; however, they require larger training datasets than convolutional neural networks. To overcome this obstacle, data-efficient ViTs were proposed, but they are typically trained using a single source of data, which overlooks the valuable knowledge that could be leveraged from other available datasets. Naïvly combining datasets from different domains can result in negative knowledge transfer (NKT), i.e., a decrease in model performance on some domains with non-negligible inter-domain heterogeneity. In this paper, we propose MDViT, the first multi-domain ViT that includes domain adapters to mitigate data-hunger and combat NKT by adaptively exploiting knowledge in multiple small data resources (domains). Further, to enhance representation learning across domains, we integrate a mutual knowledge distillation paradigm that transfers knowledge between a universal network (spanning all the domains) and auxiliary domain-specific branches. Experiments on 4 skin lesion segmentation datasets show that MDViT outperforms state-of-the-art algorithms, with superior segmentation performance and a fixed model size, at inference time, even as more domains are added. Our code is available at https://github.com/siyi-wind/MDViT.
@@ 1 Introduction
[7, 39, 22]
$\surd$
$\surd$
by adding inductive bias
$\times$
[31, 37]
$\surd$
$\surd$
by knowledge sharing
$\times$
[34]
$\surd$
$\surd$
by increasing dataset size
$\times$
[8]
$\surd$
$\surd$
by unsupervised pretraining
$\times$
[28, 35]
$\times$
$\times$
$\surd$
$\surd$
[21, 26]
$\times$
$\times$
$\surd$
$\times$
[32]
$\surd$
$\times$
$\surd$
$\times$
MDViT
$\surd$
$\surd$
by multi-domain learning
$\surd$
$\surd$
@@
Given the inter-domain heterogeneity resulting from variations in imaging protocols, scanner manufacturers, etc. [4, 21], directly mixing all the datasets for training, i.e., joint training, may improve a model’s performance on one dataset while degrading performance on other datasets with non-negligible unrelated domain-specific information, a phenomenon referred to as negative knowledge transfer (NKT) [1, 38]. A common strategy to mitigate NKT in computer vision is to introduce adapters aiding the model to adapt to different domains, i.e., multi-domain adaptive training (MAT), such as domain-specific mechanisms [21, 26, 32], and squeeze-excitation layers [35, 28] (Table 1). However, those MAT techniques were built based on CNN rather than ViT or were scalable, i.e., the models’ size at the inference time increases linearly with the number of domains.
@@ 2 Methodology
Let $\boldsymbol{X}\in\mathbb{R}^{H\times W\times 3}$ be an input RGB image and $\boldsymbol{Y}\in\{0,1\}^{H\times W}$ be its ground-truth segmentation mask. Training samples $\{(\boldsymbol{X},\boldsymbol{Y})\}$ come from $M$ datasets, each representing a domain. We aim to build and train a single ViT that performs well on all domain data and addresses the insufficiency of samples in any of the datasets. We first introduce our baseline (BASE), a ViT with hierarchical transformer blocks (Fig. 1(a)). Our proposed MDViT extends BASE with 1) a domain adapter (DA) module inside the factorized multi-head self-attention (MHSA) to adapt the model to different domains (Fig. 1(b,c)), and 2) a mutual knowledge distillation (MKD) strategy to extract more robust representations across domains (Fig. 1(d)). We present the details of MDViT in Section 2.1.
@@ 2.1 MDViT
Domain Adapter (DA): In multi-domain adaptive training, some methods build domain-specific layers in parallel with the main network [26, 21, 32]. Without adding domain-specific layers, we utilize the existing parallel structure in ViTs, i.e., MHSA, for domain adaptation. The $H$ parallel heads of MHSA mimic how humans examine the same object from different perspectives [10]. Similarly, our intuition of inserting the DA into MHSA is to enable the different heads to have varied perspectives across domains. Rather than manually designate each head to one of the domains, guided by a domain label, MDViT learns to focus on the corresponding features from different heads when encountering a domain. DA contains two steps: Attention Generation and Information Selection (Fig. 1(c)).
@@
Mutual Knowledge Distillation (MKD): Distilling knowledge from domain-specific networks has been found beneficial for universal networks to learn more robust representations [21, 40]. Moreover, mutual learning that transfers knowledge between teachers and students enables both to be optimized simultaneously [15]. To realize these benefits, we propose MKD that mutually transfers knowledge between auxiliary peers and the universal network. In Fig. 1(d), the $m$th auxiliary peer is only trained on the $m$th domain, producing output $\hat{\boldsymbol{Y}}^{m}$, whereas the universal network’s output is $\hat{\boldsymbol{Y}}$. Similar to [21], we utilize a symmetric Dice loss $L_{mkd}^{a_{m}}=Dice(\hat{\boldsymbol{Y}},\hat{\boldsymbol{Y}}^{m})$ as the knowledge distillation loss. Each peer is an expert in a certain domain, guiding the universal network to learn domain-specific information. The universal network experiences all the domains and grasps the domain-shared knowledge, which is beneficial for peer learning.
Each Auxiliary Peer is trained on a small, individual dataset specific to that peer (Fig. 1(d)). To achieve a rapid training process and prevent overfitting, particularly when working with numerous training datasets, we adapt a lightweight multilayer perception (MLP) decoder designed for ViT encoders [36] to our peers’ architecture. Specifically, multi-level features from the encoding transformer blocks (Fig. 1(a)) go through an MLP layer and an up-sample operation to unify the channel dimension and resolution to $\frac{H}{4}\times\frac{W}{4}$, which are then concatenated with the feature involving domain-shared information from the universal network’s last transformer block. Finally, we pass the fused feature to an MLP layer and do an up-sample to obtain a segmentation map.
@@ 2.2 Objective Function
Table 2: Segmentation results comparing BASE, MDViT, and SOTA methods. We report the models’ parameter count at inference time in millions (M). T means training paradigms. ${}^{\dagger}$ represents using domain-specific normalization.
@@
(millions) Dice normal-$\uparrow$ IOU normal-$\uparrow$
ISIC
DMF
SCD
PH2
avg $\pm$ std
ISIC
DMF
SCD
PH2
avg $\pm$ std
@@
BASE
27.8$\times$
ST
90.18
90.68
86.82
93.41
90.27 $\pm$ 1.16
82.82
83.22
77.64
87.84
82.88 $\pm$ 1.67
BASE
27.8
JT
89.42
89.89
92.96
94.24
91.63 $\pm$ 0.42
81.68
82.07
87.03
89.36
85.04 $\pm$ 0.64
@@
SwinUnet
41.4$\times$
ST
89.25
90.69
88.58
94.13
90.66 $\pm$ 0.87
81.51
83.25
80.40
89.00
83.54 $\pm$ 1.27
SwinUnet
41.4
JT
89.64
90.40
92.98
94.86
91.97 $\pm$ 0.30
81.98
82.80
87.08
90.33
85.55 $\pm$ 0.50
UTNet
10.0$\times$
ST
89.74
90.01
88.13
93.23
90.28 $\pm$ 0.62
82.16
82.13
79.87
87.60
82.94 $\pm$ 0.82
UTNet
10.0
JT
90.24
89.85
92.06
94.75
91.72 $\pm$ 0.63
82.92
82.00
85.66
90.17
85.19 $\pm$ 0.96
BAT
32.2$\times$
ST
90.45
90.56
90.78
94.72
91.63 $\pm$ 0.68
83.04
82.97
83.66
90.03
84.92 $\pm$ 1.01
BAT
32.2
JT
90.06
90.06
92.66
93.53
91.58 $\pm$ 0.33
82.44
82.18
86.48
88.11
84.80 $\pm$ 0.53
TransFuse
26.3$\times$
ST
90.43
91.04
91.37
94.93
91.94 $\pm$ 0.67
83.18
83.86
84.91
90.44
85.60 $\pm$ 0.95
TransFuse
26.3
JT
90.03
90.48
92.54
95.14
92.05 $\pm$ 0.36
82.56
82.97
86.50
90.85
85.72 $\pm$ 0.56
Swin UNETR
25.1$\times$
ST
90.29
90.95
91.10
94.45
91.70 $\pm$ 0.51
82.93
83.69
84.16
89.59
85.09 $\pm$ 0.79
Swin UNETR
25.1
JT
89.81
90.87
92.29
94.73
91.93 $\pm$ 0.29
82.21
83.58
86.10
90.11
85.50 $\pm$ 0.44
@@
Rundo et al.
28.2
MAT
89.43
89.46
92.62
94.68
91.55 $\pm$ 0.64
81.73
81.40
86.71
90.12
84.99 $\pm$ 0.90
Wang et al.
28.1
MAT
89.46
89.62
92.62
94.47
91.55 $\pm$ 0.54
81.79
81.59
86.71
89.76
84.96 $\pm$ 0.74
${}^{\dagger}$
27.8(.02$\times$)
MAT
90.22
90.61
93.69
95.55
92.52 $\pm$ 0.45
82.91
83.14
88.28
91.58
86.48 $\pm$ 0.74
${}^{\dagger}$
28.6(.02$\times$)
MAT
90.24
90.71
93.38
95.90
92.56 $\pm$ 0.52
82.97
83.31
88.06
92.19
86.64 $\pm$ 0.76
@@ 3 Experiments
Implementation Details: We conduct 3 training paradigms: separate (ST), joint (JT), and multi-domain adaptive training (MAT), described in Section 1, to train all the models from scratch on the skin datasets. Images are resized to $256\times 256$ and then augmented through random scaling, shifting, rotation, flipping, Gaussian noise, and brightness and contrast changes. The encoding transformer blocks’ channel dimensions are [64, 128, 320, 512] (Fig. 1(a)). We use two transformer layers in each transformer block and set the number of heads in MHSA to 8. The hidden dimensions of the CNN bridge and auxiliary peers are 1024 and 512. We deploy models on a single TITAN V GPU and train them for 200 epochs with the AdamW [23] optimizer, a batch size of 16, ensuring 4 samples from each dataset, and an initial learning rate of $1\times 10^{-4}$, which changes through a linear decay scheduler whose step size is 50 and decay factor $\gamma=0.5$.
Comparing Against BASE: In Table 2(a,b), compared with BASE in ST, BASE in JT improves the segmentation performance on small datasets (PH2 and SCD) but at the expense of diminished performance on larger datasets (ISIC and DMF). This is expected given the non-negligible inter-domain heterogeneity between skin lesion datasets, as found by Bayasi et al. [5]. The above results demonstrate that shared knowledge in related domains facilitates training a ViT on small datasets while, without a well-designed multi-domain algorithm, causing negative knowledge transfer (NKT) due to inter-domain heterogeneity, i.e., the model’s performance decreases on other datasets. Meanwhile, MDViT fits all the domains without NKT and outperforms BASE in ST by a large margin; significantly increasing Dice and IOU on SCD by 6.4% and 10.16%, showing that MDViT smartly selects valuable knowledge when given data from a certain domain. Additionally, MDViT outperforms BASE in JT across all the domains, with average improvements of 0.82% on Dice and 1.4% on IOU.
@@
Comparing Against State-of-the-Art (SOTA) Methods: We conduct experiments on SOTA data-efficient MIS ViTs and multi-domain learning methods. Previous MIS ViTs mitigated the data-hunger in one dataset by adding inductive bias, e.g., SwinUnet [7], UTNet [13], BAT [33], TransFuse [39], and Swin UNETR [30]. We implement ResNet-34 as the backbone of BAT for fair comparison (similar model size). As illustrated in Table 2(a,b,c), these SOTA models are superior to BASE in SJ. This is expected since they are designed to reduce data requirements. Nevertheless, in JT, these models also suffer from NKT: They perform better than models in ST on some datasets, like SCD, and worse on others, like ISIC. Finally, MDViT achieves the best segmentation performance in average Dice and IOU without NKT and has the best results on SCD and PH2. Fig. 2 shows MDViT’s excellent performance on ISIC and DMF and that it achieves of the closest results to ground truth on SCD and PH2. More segmentation results are presented in the supplementary material. Though BAT and TransFuse in ST have better results on some datasets like ISIC, they require extra compute resources to train $M$ models as well as an $M$-fold increase in memory requirements. The above results indicate that domain-shared knowledge is especially beneficial for training relatively small datasets such as SCD.
We employ the two fixed-size (i.e., independent of $M$) multi-domain algorithms proposed by Rundo et al. [28] and Wang et al. [35] on BASE. We set the number of parallel SE adapters in [35] to 4. In Table 2(b,d), MDViT outperforms both of them on all the domains, showing the efficacy of MDViT and that multi-domain methods built on ViTs might not perform as well as on CNNs. We also apply the domain-specific normalization [21] to BASE and MDViT to get ${}^{\dagger}$ and ${}^{\dagger}$, respectively. In Table 2(d), ${}^{\dagger}$ confronts NKT, which lowers the performance on DMF compared with BASE in ST, whereas ${}^{\dagger}$ not only addresses NKT but also outperforms ${}^{\dagger}$ on average Dice and IOU.
Table 3: Ablation studies of MDViT and experiments of DA’s plug-in capability. KD means general knowledge distillation, i.e., we only transfer knowledge from auxiliary peers to the universal network. ${}^{D}$ or ${}^{B}$ refers to using DeepLabv3’s decoder or BASE’s decoding layers as auxiliary peers.
@@
DosViT
14.6
JT
88.66
89.72
90.65
94.26
90.82 $\pm$ 0.43
80.45
81.73
83.29
89.26
83.68 $\pm$ 0.68
DosViT+DA
14.9
MAT
89.22
89.91
90.73
94.42
91.07 $\pm$ 0.32
81.28
82.00
83.44
89.57
84.07 $\pm$ 0.50
TransFuse
26.3
JT
90.03
90.48
92.54
95.14
92.05 $\pm$ 0.36
82.56
82.97
86.50
90.85
85.72 $\pm$ 0.56
TransFuse+DA
26.9
MAT
90.13
90.47
93.62
95.21
92.36 $\pm$ 0.38
82.80
82.94
88.16
90.97
86.22 $\pm$ 0.64
@@
BASE
27.8
JT
89.42
89.89
92.96
94.24
91.63 $\pm$ 0.42
81.68
82.07
87.03
89.36
85.04 $\pm$ 0.64
BASE+DA
28.5
MAT
89.96
90.66
93.36
95.46
92.36 $\pm$ 0.51
82.52
83.24
87.98
91.43
86.29 $\pm$ 0.72
BASE+MKD
27.8
JT
89.27
89.53
92.66
94.83
91.57 $\pm$ 0.53
81.45
81.49
86.81
90.42
85.04 $\pm$ 0.74
BASE+DA+KD
28.5
MAT
90.03
90.59
93.26
95.63
92.38 $\pm$ 0.39
82.67
83.12
87.85
91.72
86.34 $\pm$ 0.51
@@
[…]
28.5
MAT
89.64
90.25
92.24
95.36
91.87 $\pm$ 0.45
82.10
82.55
86.12
91.24
85.50 $\pm$ 0.67
[…]
28.5
MAT
90.03
90.73
92.72
95.32
92.20 $\pm$ 0.50
82.66
83.35
87.01
91.17
86.05 $\pm$ 0.70
MDViT
28.5
MAT
90.29
90.78
93.22
95.53
92.45 $\pm$ 0.65
82.99
83.41
87.80
91.57
86.44 $\pm$ 0.94
Ablation Studies and Plug-in Capability of DA: We conduct ablation studies to demonstrate the efficacy of DA, MKD, and auxiliary peers. Table 3(b) reveals that using one-direction knowledge distillation (KD) or either of the critical components in MDViT, i.e., DA or MKD, but not together, could not achieve the best results. Table 3(c) exemplifies that, for building the auxiliary peers, our proposed MLP architecture is more effective and has fewer parameters ($1.6$M) than DeepLabv3’s decoder [9] ($4.7$M) or BASE’s decoding layers (10.8M). Finally, we incorporate DA into two ViTs: TransFuse and DosViT (the latter includes the earliest ViT encoder [12] and a DeepLabv3’s decoder). As shown in Table 3(a,b), DA can be used in various ViTs but is more advantageous in MDViT with more transformer blocks in the encoding and decoding process.

### v3

@@ Abstract
Despite its clinical utility, medical image segmentation (MIS) remains a daunting task due to images’ inherent complexity and variability. Vision transformers (ViTs) have recently emerged as a promising solution to improve MIS; however, they require larger training datasets than convolutional neural networks. To overcome this obstacle, data-efficient ViTs were proposed, but they are typically trained using a single source of data, which overlooks the valuable knowledge that could be leveraged from other available datasets. Naïvly combining datasets from different domains can result in negative knowledge transfer (NKT), i.e., a decrease in model performance on some domains with non-negligible inter-domain heterogeneity. In this paper, we propose MDViT, the first multi-domain ViT that includes domain adapters to mitigate data-hunger and combat NKT by adaptively exploiting knowledge in multiple small data resources (domains). Further, to enhance representation learning across domains, we integrate a mutual knowledge distillation paradigm that transfers knowledge between a universal network (spanning all the domains) and auxiliary domain-specific network branches. Experiments on 4 skin lesion segmentation datasets show that MDViT outperforms state-of-the-art algorithms, with superior segmentation performance and a fixed model size, at inference time, even as more domains are added. Our code is available at https://github.com/siyi-wind/MDViT.
@@ 1 Introduction
[7, 39, 22] $\surd$ $\surd$ by adding inductive bias $\times$ -
[31, 37] $\surd$ $\surd$ by knowledge sharing $\times$ -
[34] $\surd$ $\surd$ by increasing dataset size $\times$ -
[8] $\surd$ $\surd$ by unsupervised pretraining $\times$ -
[28, 35] $\times$ $\times$ $\surd$ $\surd$
[21, 26] $\times$ $\times$ $\surd$ $\times$
[32] $\surd$ $\times$ $\surd$ $\times$
MDViT $\surd$ $\surd$ by multi-domain learning $\surd$ $\surd$
@@
Given the inter-domain heterogeneity resulting from variations in imaging protocols, scanner manufacturers, etc. [4, 21], directly mixing all the datasets for training, i.e., joint training, may improve a model’s performance on one dataset while degrading performance on other datasets with non-negligible unrelated domain-specific information, a phenomenon referred to as negative knowledge transfer (NKT) [1, 38]. A common strategy to mitigate NKT in computer vision is to introduce adapters aiding the model to adapt to different domains, i.e., multi-domain adaptive training (MAT), such as domain-specific mechanisms [21, 26, 32], and squeeze-excitation layers [35, 28] (Table 1). However, those MAT techniques are built based on CNN rather than ViT or are scalable, i.e., the models’ size at the inference time increases linearly with the number of domains.
@@ 2 Methodology
Let $\boldsymbol{X}\in\mathbb{R}^{H\times W\times 3}$ be an input RGB image and $\boldsymbol{Y}\in\{0,1\}^{H\times W}$ be its ground-truth segmentation mask. Training samples $\{(\boldsymbol{X},\boldsymbol{Y})\}$ come from $M$ datasets, each representing a domain. We aim to build and train a single ViT that performs well on all domain data and addresses the insufficiency of samples in any of the datasets. We first introduce our baseline (BASE), a ViT with hierarchical transformer blocks (Fig. 1-a). Our proposed MDViT extends BASE with 1) a domain adapter (DA) module inside the factorized multi-head self-attention (MHSA) to adapt the model to different domains (Fig. 1-b,c), and 2) a mutual knowledge distillation (MKD) strategy to extract more robust representations across domains (Fig. 1-d). We present the details of MDViT in Section 2.1.
@@ 2.1 MDViT
Domain Adapter (DA): In multi-domain adaptive training, some methods build domain-specific layers in parallel with the main network [26, 21, 32]. Without adding domain-specific layers, we utilize the existing parallel structure in ViTs, i.e., MHSA, for domain adaptation. The $H$ parallel heads of MHSA mimic how humans examine the same object from different perspectives [10]. Similarly, our intuition of inserting the DA into MHSA is to enable the different heads to have varied perspectives across domains. Rather than manually designate each head to one of the domains, guided by a domain label, MDViT learns to focus on the corresponding features from different heads when encountering a domain. DA contains two steps: Attention Generation and Information Selection (Fig. 1-c).
@@
Mutual Knowledge Distillation (MKD): Distilling knowledge from domain-specific networks has been found beneficial for universal networks to learn more robust representations [21, 40]. Moreover, mutual learning that transfers knowledge between teachers and students enables both to be optimized simultaneously [15]. To realize these benefits, we propose MKD that mutually transfers knowledge between auxiliary peers and the universal network. In Fig. 1-d, the $m$th auxiliary peer is only trained on the $m$th domain, producing output $\hat{\boldsymbol{Y}}^{m}$, whereas the universal network’s output is $\hat{\boldsymbol{Y}}$. Similar to [21], we utilize a symmetric Dice loss $L_{mkd}^{a_{m}}=Dice(\hat{\boldsymbol{Y}},\hat{\boldsymbol{Y}}^{m})$ as the knowledge distillation loss. Each peer is an expert in a certain domain, guiding the universal network to learn domain-specific information. The universal network experiences all the domains and grasps the domain-shared knowledge, which is beneficial for peer learning.
Each Auxiliary Peer is trained on a small, individual dataset specific to that peer (Fig. 1-d). To achieve a rapid training process and prevent overfitting, particularly when working with numerous training datasets, we adapt a lightweight multilayer perception (MLP) decoder designed for ViT encoders [36] to our peers’ architecture. Specifically, multi-level features from the encoding transformer blocks (Fig. 1-a) go through an MLP layer and an up-sample operation to unify the channel dimension and resolution to $\frac{H}{4}\times\frac{W}{4}$, which are then concatenated with the feature involving domain-shared information from the universal network’s last transformer block. Finally, we pass the fused feature to an MLP layer and do an up-sample to obtain a segmentation map.
@@ 2.2 Objective Function
Table 2: Segmentation results comparing BASE, MDViT, and SOTA methods. We report the models’ parameter count at inference time in millions (M). T means training paradigms. represents using domain-specific normalization.
@@
(millions) Dice $\uparrow$ IOU $\uparrow$
(M) ISIC DMF SCD PH2 avg $\pm$ std ISIC DMF SCD PH2 avg $\pm$ std
@@
BASE 27.8$\times$ ST 90.18 90.68 86.82 93.41 90.27 $\pm$ 1.16 82.82 83.22 77.64 87.84 82.88 $\pm$ 1.67
BASE 27.8 JT 89.42 89.89 92.96 94.24 91.63 $\pm$ 0.42 81.68 82.07 87.03 89.36 85.04 $\pm$ 0.64
@@
SwinUnet 41.4$\times$ ST 89.25 90.69 88.58 94.13 90.66 $\pm$ 0.87 81.51 83.25 80.40 89.00 83.54 $\pm$ 1.27
SwinUnet 41.4 JT 89.64 90.40 92.98 94.86 91.97 $\pm$ 0.30 81.98 82.80 87.08 90.33 85.55 $\pm$ 0.50
UTNet 10.0$\times$ ST 89.74 90.01 88.13 93.23 90.28 $\pm$ 0.62 82.16 82.13 79.87 87.60 82.94 $\pm$ 0.82
UTNet 10.0 JT 90.24 89.85 92.06 94.75 91.72 $\pm$ 0.63 82.92 82.00 85.66 90.17 85.19 $\pm$ 0.96
BAT 32.2$\times$ ST 90.45 90.56 90.78 94.72 91.63 $\pm$ 0.68 83.04 82.97 83.66 90.03 84.92 $\pm$ 1.01
BAT 32.2 JT 90.06 90.06 92.66 93.53 91.58 $\pm$ 0.33 82.44 82.18 86.48 88.11 84.80 $\pm$ 0.53
TransFuse 26.3$\times$ ST 90.43 91.04 91.37 94.93 91.94 $\pm$ 0.67 83.18 83.86 84.91 90.44 85.60 $\pm$ 0.95
TransFuse 26.3 JT 90.03 90.48 92.54 95.14 92.05 $\pm$ 0.36 82.56 82.97 86.50 90.85 85.72 $\pm$ 0.56
Swin UNETR 25.1$\times$ ST 90.29 90.95 91.10 94.45 91.70 $\pm$ 0.51 82.93 83.69 84.16 89.59 85.09 $\pm$ 0.79
Swin UNETR 25.1 JT 89.81 90.87 92.29 94.73 91.93 $\pm$ 0.29 82.21 83.58 86.10 90.11 85.50 $\pm$ 0.44
@@
Rundo et al. 28.2 MAT 89.43 89.46 92.62 94.68 91.55 $\pm$ 0.64 81.73 81.40 86.71 90.12 84.99 $\pm$ 0.90
Wang et al. 28.1 MAT 89.46 89.62 92.62 94.47 91.55 $\pm$ 0.54 81.79 81.59 86.71 89.76 84.96 $\pm$ 0.74
BASE 27.8(.02$\times$) MAT 90.22 90.61 93.69 95.55 92.52 $\pm$ 0.45 82.91 83.14 88.28 91.58 86.48 $\pm$ 0.74
MDViT 28.6(.02$\times$) MAT 90.24 90.71 93.38 95.90 92.56 $\pm$ 0.52 82.97 83.31 88.06 92.19 86.64 $\pm$ 0.76
@@ 3 Experiments
Implementation Details: We conduct 3 training paradigms: separate (ST), joint (JT), and multi-domain adaptive training (MAT), described in Section 1, to train all the models from scratch on the skin datasets. Images are resized to $256\times 256$ and then augmented through random scaling, shifting, rotation, flipping, Gaussian noise, and brightness and contrast changes. The encoding transformer blocks’ channel dimensions are [64, 128, 320, 512] (Fig. 1-a). We use two transformer layers in each transformer block and set the number of heads in MHSA to 8. The hidden dimensions of the CNN bridge and auxiliary peers are 1024 and 512. We deploy models on a single TITAN V GPU and train them for 200 epochs with the AdamW [23] optimizer, a batch size of 16, ensuring 4 samples from each dataset, and an initial learning rate of $1\times 10^{-4}$, which changes through a linear decay scheduler whose step size is 50 and decay factor $\gamma=0.5$.
Comparing Against BASE: In Table 2-a,b, compared with BASE in ST, BASE in JT improves the segmentation performance on small datasets (PH2 and SCD) but at the expense of diminished performance on larger datasets (ISIC and DMF). This is expected given the non-negligible inter-domain heterogeneity between skin lesion datasets, as found by Bayasi et al. [5]. The above results demonstrate that shared knowledge in related domains facilitates training a ViT on small datasets while, without a well-designed multi-domain algorithm, causing negative knowledge transfer (NKT) due to inter-domain heterogeneity, i.e., the model’s performance decreases on other datasets. Meanwhile, MDViT fits all the domains without NKT and outperforms BASE in ST by a large margin; significantly increasing Dice and IOU on SCD by 6.4% and 10.16%, showing that MDViT smartly selects valuable knowledge when given data from a certain domain. Additionally, MDViT outperforms BASE in JT across all the domains, with average improvements of 0.82% on Dice and 1.4% on IOU.
@@
Comparing Against State-of-the-Art (SOTA) Methods: We conduct experiments on SOTA data-efficient MIS ViTs and multi-domain learning methods. Previous MIS ViTs mitigated the data-hunger in one dataset by adding inductive bias, e.g., SwinUnet [7], UTNet [13], BAT [33], TransFuse [39], and Swin UNETR [30]. We implement ResNet-34 as the backbone of BAT for fair comparison (similar model size). As illustrated in Table 2-a,b,c, these SOTA models are superior to BASE in SJ. This is expected since they are designed to reduce data requirements. Nevertheless, in JT, these models also suffer from NKT: They perform better than models in ST on some datasets, like SCD, and worse on others, like ISIC. Finally, MDViT achieves the best segmentation performance in average Dice and IOU without NKT and has the best results on SCD and PH2. Fig. 2 shows MDViT’s excellent performance on ISIC and DMF and that it achieves the closest results to ground truth on SCD and PH2. More segmentation results are presented in the supplementary material. Though BAT and TransFuse in ST have better results on some datasets like ISIC, they require extra compute resources to train $M$ models as well as an $M$-fold increase in memory requirements. The above results indicate that domain-shared knowledge is especially beneficial for training relatively small datasets such as SCD.
We employ the two fixed-size (i.e., independent of $M$) multi-domain algorithms proposed by Rundo et al. [28] and Wang et al. [35] on BASE. We set the number of parallel SE adapters in [35] to 4. In Table 2-b,d, MDViT outperforms both of them on all the domains, showing the efficacy of MDViT and that multi-domain methods built on ViTs might not perform as well as on CNNs. We also apply the domain-specific normalization [21] to BASE and MDViT to get BASE and MDViT, respectively. In Table 2-d, BASE confronts NKT, which lowers the performance on DMF compared with BASE in ST, whereas MDViT not only addresses NKT but also outperforms BASE on average Dice and IOU.
Table 3: Ablation studies of MDViT and experiments of DA’s plug-in capability. KD means general knowledge distillation, i.e., we only transfer knowledge from auxiliary peers to the universal network. D or B refers to using DeepLabv3’s decoder or BASE’s decoding layers as auxiliary peers.
@@
DosViT 14.6 JT 88.66 89.72 90.65 94.26 90.82 $\pm$ 0.43 80.45 81.73 83.29 89.26 83.68 $\pm$ 0.68
DosViT+DA 14.9 MAT 89.22 89.91 90.73 94.42 91.07 $\pm$ 0.32 81.28 82.00 83.44 89.57 84.07 $\pm$ 0.50
TransFuse 26.3 JT 90.03 90.48 92.54 95.14 92.05 $\pm$ 0.36 82.56 82.97 86.50 90.85 85.72 $\pm$ 0.56
TransFuse+DA 26.9 MAT 90.13 90.47 93.62 95.21 92.36 $\pm$ 0.38 82.80 82.94 88.16 90.97 86.22 $\pm$ 0.64
@@
BASE 27.8 JT 89.42 89.89 92.96 94.24 91.63 $\pm$ 0.42 81.68 82.07 87.03 89.36 85.04 $\pm$ 0.64
BASE+DA 28.5 MAT 89.96 90.66 93.36 95.46 92.36 $\pm$ 0.51 82.52 83.24 87.98 91.43 86.29 $\pm$ 0.72
BASE+MKD 27.8 JT 89.27 89.53 92.66 94.83 91.57 $\pm$ 0.53 81.45 81.49 86.81 90.42 85.04 $\pm$ 0.74
BASE+DA+KD 28.5 MAT 90.03 90.59 93.26 95.63 92.38 $\pm$ 0.39 82.67 83.12 87.85 91.72 86.34 $\pm$ 0.51
@@
MDViTD 28.5 MAT 89.64 90.25 92.24 95.36 91.87 $\pm$ 0.45 82.10 82.55 86.12 91.24 85.50 $\pm$ 0.67
MDViTB 28.5 MAT 90.03 90.73 92.72 95.32 92.20 $\pm$ 0.50 82.66 83.35 87.01 91.17 86.05 $\pm$ 0.70
MDViT 28.5 MAT 90.29 90.78 93.22 95.53 92.45 $\pm$ 0.65 82.99 83.41 87.80 91.57 86.44 $\pm$ 0.94
Ablation Studies and Plug-in Capability of DA: We conduct ablation studies to demonstrate the efficacy of DA, MKD, and auxiliary peers. Table 3-b reveals that using one-direction knowledge distillation (KD) or either of the critical components in MDViT, i.e., DA or MKD, but not together, could not achieve the best results. Table 3-c exemplifies that, for building the auxiliary peers, our proposed MLP architecture is more effective and has fewer parameters ($1.6$M) than DeepLabv3’s decoder [9] ($4.7$M) or BASE’s decoding layers (10.8M). Finally, we incorporate DA into two ViTs: TransFuse and DosViT (the latter includes the earliest ViT encoder [12] and a DeepLabv3’s decoder). As shown in Table 3-a,b, DA can be used in various ViTs but is more advantageous in MDViT with more transformer blocks in the encoding and decoding process.